A man borrows $360. If he pays it back in 12 monthly installments of$31.50, what is his interest rate?
A)1.5%
B)4.5%
C)10%
D)5%
E)7.5%

Answer by jim_thompson5910(35256)

"he pays it back in 12 monthly installments of$31.50" means he pays a total of 12*31.50 = 378 dollars.

Subtract 360 from this: 378 - 360 = 18

So he pays a total of $18 in interest. This means I = 18

The principle is 360. So P = 360

The amount of time is one year, so t = 1


I = Prt

18 = 360*r*1

18 = 360r

18/360 = r

0.05 = r

r = 0.05

So the interest rate is 5%